### v2

# Propagation Delays of Ultra-High-Energy Cosmic Rays from Active Galactic Nuclei

Abstract
We investigate the time delay incurred during ultra-high energy cosmic ray (UHECR) propagation over cosmological distances and its potential impact on the correlation between UHECR directions of arrival and sources such as Active Galactic Nuclei (AGNs), the UHECR chemical composition, and extragalactic magnetic field constraints. We propagate particles in different magnetic field configurations, spanning over an extended range of particle Larmor radii and magnetic field coherence lengths, also including attenuation losses. We find that UHECR delays could easily be comparable to (and longer than) AGN duty cycles, effectively erasing the correlation between known AGNs and UHECR anisotropies. We finally consider how strong constraints on the chemical composition of the heaviest UHECRs could enable a better characterization of extragalactic magnetic fields.
keywords:
Ultra-high-energy cosmic rays , Active Galactic Nuclei , Cosmic Rays , Particle propagation
journal: Astroparticle Physics
[…]
[label1]organization=Joint Space-Science Institute, University of Maryland, city=College Park, postcode=20742, state=MD, country=USA
[…]
1 Introduction
Pinpointing the sources of the Ultra-High-Energy Cosmic Rays (UHECRs) may involve finding a correlation between their time/directions of arrival and the distribution of potential sources. The correlation with transient or short-lived sources, such as gamma-ray bursts and newly-born millisecond neutron stars, is already washed away by the delay induced by propagation in our Galaxy, which is likely of the order of $\sim$1000 yr (e.g., Miralda-Escudé and Waxman, 1996; Murase, 2009; Kalli, S. et al., 2011; Takami and Murase, 2012; Harari et al., 2021; van Vliet et al., 2021). For Active Galactic Nuclei (AGNs), correlation studies have mainly focused on flaring phases that would coincide with UHECR acceleration (e.g., Biermann and Strittmatter, 1987; Dermer et al., 2009; Murase et al., 2012; Farrar and Gruzinov, 2009), but little attention has been paid to analyzing propagation delays that could be compared with their much longer duty cycles.
In this work, we consider a large range of possible properties of the magnetic fields between sources and Earth to examine potential UHECR delays during propagation, and their impact on detected UHECR chemical composition. We find that UHECR propagation can introduce delays comparable to AGN duty cycles, rendering statistical correlations and anisotropy studies involving AGNs quite challenging; moreover, these conclusions likely extend to starburst galaxies, since the periods of enhanced star formation could be comparable to AGN duty cycles, for instance when fed by major galaxy mergers. Star formation can be quenched either by AGN or stellar feedback; in the former case there is a direct physical connection between the two duty cycles, in the latter quenching should kick in when massive stars go supernovae, which is still in the Myr ballpark. Within the order of magnitude uncertainty that applies to AGN duty cycles, we conclude that the arguments in this paper should apply to scenarios in which UHECR sources are either starburst galaxies or AGNs.
1.1 AGN duty Cycles
[…]
1.2 Magnetic Fields and UHECR Propagation
Constraining the properties of magnetic fields between UHECR sources and Earth is of paramount importance for their propagation and for evaluating the effects of propagation delays. In this regard, the effects of the galactic magnetic field (GMF), intergalactic magnetic field (IGMF), the intra-cluster medium (ICM) $B$-fields, and even $B$-fields in voids should be important.
UHECRs can experience important deflections because of the GMF (e.g., Takami et al., 2012; Jansson and Farrar, 2012; Alves Batista et al., 2016; Unger and Farrar, 2017, 2024), especially because GMFs can easily reach $B\sim 1$-$10\mu$G. In this regard, the effects of the GMF toroidal component (Sun et al., 2008; Pshirkov et al., 2011), and poloidal and random components (Jansson and Farrar, 2012; Farrar and Sutherland, 2019) have been found to potentially cause significant deflections. Generally speaking, the structure of the GMF remains poorly constrained, but its effects on UHECR deflection should be inevitable, with delays of the order of $\sim 10^{3}$yr for our Galaxy (e.g., Murase, 2009).
As for the extragalactic components, numerous studies have attempted to constrain the properties of intergalactic magnetic fields (IGMFs) (e.g., Neronov and Vovk, 2010; Tavecchio et al., 2010, 2011; Huan et al., 2011; Taylor et al., 2011; Finke et al., 2015) and their corresponding coherence lengths (e.g., Vovk et al., 2012). However, the scarcity of observational data and the difficulties associated with measurements of $B$-fields (Alves Batista and Saveliev, 2021) have made this challenging. Moreover, in the context of quasars, outflows such as jets can inject quite substantial fields (e.g., Kulsrud and Zweibel, 2008; Vallée, 2011; Ryu et al., 2012) in the IGM (Furlanetto and Loeb, 2001).
As for the ICM, it may also be populated with non-negligible fields, especially that filaments of galaxy clusters are found to have $B$-fields reaching $B\sim 0.1$-$10$nG (e.g., Vazza et al., 2017), with maximum values based on magnetic draping reaching a few $\mu$G (Pfrommer and Dursi, 2010). Faraday rotation observations also suggest that $B$-fields can reach $\sim\mu$G levels (e.g., Clarke et al., 2001; Carilli and Taylor, 2002; Govoni et al., 2004). The local galaxy supercluster, on the other hand, has been found to have a statistically significant “Faraday screen” acting on radio waves, which would return a field as large as $B\sim 0.3\mu$G in the local Universe (Vallée, 2002). Similarly, the plane of the Coma supercluster can reach 0.5-1.5 $\mu$G, declining to about 0.03-0.5 $\mu$G on $\gtrsim 1$ Mpc levels (Brunetti et al., 2001). Other work demonstrated that radio halos in the ICM are diffuse Mpc-scale synchrotron sources produced by GeV electrons interacting with $\mu$G-scale magnetic fields (Cassano et al., 2010; Cuciti et al., 2022). Finally, simulated $B$-fields in the local supercluster find that they can reach $\sim 10$nG (Ryu et al., 2008).
It is worth mentioning that regions with indications of UHECR excesses have been reported to intersect the supergalactic plane (e.g., Aab et al., 2018; Abdul Halim et al., 2023; Kim et al., 2023; The Pierre Auger Collaboration et al., 2024; Globus et al., 2019), and that the UHECR dipole anisotropy could be associated with the orientation of the Galaxy relative to the supergalactic plane (Globus and Piran, 2017; Globus et al., 2019; Ding et al., 2021; Bister and Farrar, 2024). Other studies suggested that cluster $B$-fields were $\sim 3\mu$G based on synchrotron cooling of populations of secondary electrons inferred from the correlation between the radio and thermal X-ray luminosity of radio emitting galaxy clusters (Kushnir et al., 2009). Within the context of UHECR propagation over cosmological baselines, it has been proposed that $B$-fields in voids are the most decisive component (Alves Batista and Saveliev, 2021), and an upper limit for their values could be set at $\lesssim 1$nG (e.g., Blasi et al., 1999; Kronberg et al., 2007; Planck Collaboration et al., 2016; Pshirkov et al., 2016).
Despite the substantial body of work that has attempted to constrain the properties of $B$-fields in galaxies, galaxy clusters and beyond, their basic properties in regions that separate us from UHECR sources remain poorly known.
Previous studies have investigated the role of propagation delays in shaping the observed UHECRS flux and features. Matthews et al. (2018) examined whether past activity in nearby AGNs, Centaurus A and Fornax A, could explain the present-day UHECR flux, accounting for a possible efficient past AGN activity and the time delays accumulated during propagation. Eichmann et al. (2022); Eichmann and Kachelrieß (2023) analyzed how source lifetimes and magnetic deflections affect spectral hardening and composition evolution, highlighting the importance of extended propagation times in shaping observed spectra. In the context of the Local Supercluster, Mollerach and Roulet (2019, 2022) explored the effects of CR diffusion including finite source ages, emphasizing their influence on spectral features, time delays, and anisotropy. Adding to this picture, Fargion et al. (2018) estimated that AGN flare episodes are unlikely to correlate with UHECR arrival times, given potential propagation delays reaching up to $10^{6}$ years. In general, even relatively small delays ($\lesssim 1$Myr) make associations of transient activity and UHECR detections very challenging.
Collectively, these works have deepened our understanding of how magnetic fields, source characteristics, and propagation effects modulate UHECR observables.
Our work extends these efforts by developing a generalized framework for propagation delays that, crucially, incorporates the effects of the most updated theory of CR scattering and can be applied to several different environments: inside/around the sources, in the host galaxy and cluster, and on cosmological distances. In addition, we leverage UHECR chemical composition to place tighter constraints on the properties of extragalactic magnetic fields. Specifically, we attempt to bracket our ignorance by propagating particles —without assuming diffusion a priori— in a large range of $B$-field strengths, spectra, and coherence lengths, also considering the effects of particle losses due to photomeson and photodisintegration interactions with the cosmic microwave background (CMB), cosmic infrared background (CIB), and cosmic optical background (COB) (Dole et al., 2006). Our goals are to compare the ensuing delays with expected AGN duty cycles to estimate: i) the significance of UHECR delay and deflection angles, and ii) the effects of such delays on losses and their repercussions on the detected UHECR chemical composition.
2 Particle propagation
We consider a simple setup where test particles are propagated with a Boris pusher (e.g., Birdsall and Langdon, 1991) with different initial pitch angles $\mu$ in a prescribed magnetic field $\mathbf{B}$. A rigidity-dependent time step is chosen to ensure proper resolution of particle gyrations, with more than five steps per gyroperiod. While the magnitude of the magnetic field, $B$, remains constant during particle propagation, a turbulent component is added to alter its direction. The field direction changes by an angle $\theta$, following a random walk every coherence length $\ell_{\rm c}$, which characterizes the local magnetic turbulence. The angle $\theta$ is drawn from a Gaussian distribution centered at zero with a standard deviation of $\pi/2$, simulating sharp, strong magnetic field bends ($\delta B/B\sim 1$). The propagated particles are representative of different species, since propagation is dependent on rigidity only (except for losses).
Particle transport has been extensively studied in turbulent boxes with magnetohydrodynamical (MHD) simulations (e.g., Beresnyak et al., 2011; Cohet and Marcowith, 2016; Dundovic et al., 2020), but this approach is unfeasible for the vast range of scales involved in UHECR propagation. Instead, frameworks where particle diffusion is simulated on synthetic fields are routinely adopted to study UHECR transport (e.g., CRPropa, Alves Batista et al., 2016). Our approach, which does not assume diffusive transport but only the random walk of the magnetic field lines, allows us to explore a wide range of Larmor radii $\mathcal{R}$ much smaller/larger than the scattering mean free path (not possible in a diffusive approximation), the field coherence length $\ell_{\rm c}$, and distance to the source. Though it ignores the actual cosmological distribution of matter in the Universe, our quite simple framework is sufficient to investigate the effects of typical strengths and coherence lengths on the propagation delays, obtained by comparing the UHECR travel time $t$ with the ballistic time $t_{b}$ from the source, i.e., the photon (or neutrino) travel time.
We now consider different transport regimes, depending on whether $\mathcal{R}\gtrless l_{c}$.
Figure 1: Time delay maps of particle propagation as a function of the particle Larmor radius $\mathcal{R}$ and coherence length $\ell_{\rm c}$ normalized to the source distance $D$. The color bar denotes the ratio of the time delay $t_{d}=t-t_{b}$ and the photon propagation time $t_{b}$. The solid grey line characterizes the boundary $\mathcal{R}=\ell_{\rm c}$. Particle Larmor radii are also plotted for the limiting case $\lambda=D$ (dashed black for $\mathcal{R}>\ell_{\rm c}$ and dotted line for $\mathcal{R}<\ell_{\rm c}$). The different maps show the impact of the $B$-field setup and initial pitch angle. Particle losses are not accounted for in these maps.
2.1 Regime with $\mathcal{R}>\ell_{\rm c}$
This regime is particularly interesting for the transport of high-energy particles, as in small B fields Larmor radii could easily exceed coherence length scales. For $\mathcal{R}\geq\ell_{\rm c}$, particles are scattered on scales smaller than their Larmor radii resulting in a scattering mean free path $\lambda$,
$\lambda\sim\ell_{\rm c}^{1-\delta}\mathcal{R}^{\delta}$ (1)
with $\delta=2$ (e.g., Shalchi and Dosch, 2009; Plotnikov et al., 2011, 2013). The properties of this regime are reproduced automatically by our Boris-transport framework, as particles are propagated self-consistently, and only the largest magnetic bends, i.e., on $\ell_{\rm c}$-scales, should have an impact on particle propagation. While testing transport in this regime, we find good agreement with the trends shown in the Monte Carlo simulations from Casse et al. (2001) and Plotnikov et al. (2011), i.e., the ratio of perpendicular to parallel diffusion falls as $(\mathcal{R}/D)^{-2}$ as expected.
2.2 Regime with $\mathcal{R}<\ell_{\rm c}$
For $\mathcal{R}/\ell_{\rm c}\ll 1$, the mean free path in turbulence with $\delta B/B\sim 1$ can also be expressed via Eq. 1, where $\delta$ is dependent on the phenomenology dictating the spectrum of turbulence (e.g., Ng et al., 2010; Treumann et al., 2015). If scattering occurs via wave-particle resonance, the quasi-linear theory (e.g., Parker, 1965; Jokipii, 1966; Forman and Gleeson, 1975; Chapman et al., 1990), returns $\delta=1/3$ in the Kolmogorov regime (Kolmogorov, 1941; Stawarz and Petrosian, 2008) and $\delta=1/2$ in the Iroshnikov-Kraichnan regime (Iroshnikov, 1964).
Recent advances in particle transport in turbulence have demonstrated that scattering on intermittent structures (magnetic field reversals at scales of the order of $\mathcal{R}\leq\ell_{\rm c}$) leads to a mean free path that can also be expressed based on Eq. 1 (Lemoine, 2023; Kempski et al., 2023), where $\delta\simeq 0.3$. In principle, $\delta$ may depend on the properties of intermittency (Lemoine, 2023), though a general theory has not been put forward, yet.
Since we only allow for magnetic structures at the coherence length scale, $\ell_{\rm c}$, our calculations do not immediately capture the effects of resonant scattering or intermittency that could control the transport at smaller scales; if left unaddressed, this limitation would lead to underestimate the transport delays for particles with $\mathcal{R}\ll\ell_{\rm c}$. To correct for this, we incorporate the influence of intermittency and resonant scattering on small scales by augmenting the propagation distance by a factor $(\ell_{\rm c}/\mathcal{R})^{\delta}$, corresponding to the fact that for $\mathcal{R}<\ell_{\rm c}$ a particle traveling a distance of $\sim\ell_{\rm c}$ generally undergoes $\ell_{\rm c}/\lambda=(\ell_{\rm c}/\mathcal{R})^{\delta}$ scattering. A posteriori, we find that varying $\delta\in[0.2,0.6]$ has a negligible impact on our conclusions about the propagation delay of UHECRs.
2.3 Particle Losses
We consider the most important attenuation mechanisms, i.e., photomeson ($p\gamma$) and photodisintegration ($A\gamma$) interactions due to collisions with the cosmic microwave, infrared, and optical backgrounds (CMB, CIB, and COB, respectively). We keep track of the evolution of the particles’ atomic mass $A$ and charge $Z$ during propagation. For the heavier UHECRs, photodisintegration is the dominant loss process.
At every time step, an interaction probability is calculated for all species, along with a photodisintegration probability for heavier nuclei based on the Giant Dipole Resonance (GDR) cross section (e.g., Tamii et al., 2022). In a photodisintegration event, a photon-absorbing nucleus changes to another specie by releasing either a neutron or a proton, based on an interaction mean free path $\lambda_{A}$, as described, e.g., in Mbarek et al. (2023). Alpha particles can also be produced through photodisintegration, further depleting heavy nuclei of their nucleons. However, this channel is generally subdominant, and its contribution varies with the nuclear species involved. For example, the photodisintegration rate of oxygen into alpha particles accounts for only a few percent of its rate into protons (Gorbunov and Osipova, 1963). While this effect is element-dependent and could influence photodisintegration outcomes, its full impact lies beyond the scope of this work.
It is worth noting that the threshold energy for alpha production can, in some cases, be lower than for neutron or proton emission (Stecker and Salamon, 1999). However, the energy dependence of the corresponding cross section tends to be broad (e.g., Odsuren et al., 2015), and modeling uncertainties—such as deviations in resonance—can further complicate predictions (Kido et al., 2023).
A comprehensive treatment of particle propagation would ideally account for the effects of photo-nuclear interactions on particle rigidity, as these interactions modify the particle’s energy and composition. Photodisintegration alters rigidity depending on the nucleon lost: losing a neutron reduces the rigidity from $\propto A/Z$ to $\propto(A-1)/Z$, while losing a proton increases it to $\propto(A-1)/(Z-1)$. For heavy nuclei (with $A\gg 1)$ like iron, the loss of a single nucleon produces only a small change in rigidity and has minimal impact. In contrast, for lighter nuclei the effect is more significant. For instance, in nitrogen, losing a proton increases the rigidity by about 10%, while losing a neutron decreases it by a similar amount. This results in a modest broadening of the rigidity distribution, centered around the initial value.
While our current results do not explicitly incorporate this effect, its influence can be estimated by summing the delays accumulated during each photodisintegration mean free path, treating the rigidity as constant between interactions and updating it after each event. Given that the overall change in rigidity remains small and symmetric around the initial value if interactions with protons and neutrons are equally probable, we do not expect this effect to substantially impact our conclusions.
3 Results
Let us consider a source at fixed distance $D$ and different combinations of $\mathcal{R}$ and $\ell_{\rm c}$, which induce a delay in particle propagation $t_{d}=t-t_{b}$, where $t$ is the total propagation time and $t_{b}$ is the ballistic travel time. Figure 1 shows the normalized time delay $t_{d}/t_{b}$, as a function of $\mathcal{R}$ and $\ell_{\rm c}$ normalized to $D$. Note that, since propagation only cares about $\mathcal{R}=E/(eZB)$, fixing the particle energy $E$ (or its rigidity, $\rho=E/Z$, where $Z$ is the charge) and choosing a value of the magnetic field are interchangeable. Particles are injected with an initial pitch angle $\mu$ to assess its effect on propagation.
In Figure 1 we overplot the scaling of $\mathcal{R}$ with $\ell_{\rm c}$ using the definition of $\lambda$ from §2.1 and §2.2, where $\mathcal{R}\equiv\lambda^{1/\delta}\ell_{\rm c}^{1-1/\delta}$ for the limiting case $\lambda=D$. Here, $\delta$ depends on the regime: $\delta\simeq 0.3$ for $\mathcal{R}\ll\ell_{\rm c}$, and $\delta=2$ for $\mathcal{R}\gg\ell_{\rm c}$. The panels in Figure 1 differ by the particle’s initial pitch angle $\mu$, which is only relevant when $\ell_{\rm c}>D$, where scattering is minimal but delay may still be non-negligible. In the following, we examine different regions of the maps with increasing delays (marked with letters a to d), delimited by gray lines in Figure 1:
Region a
In this region ($\mathcal{R}>D$) delays are minimal and associated with a slight bending from ballistic trajectories, such that $t_{d}/t_{b}\ll 1$.
Region b
In this case $\ell_{\rm c}>D$ and the B field remains coherent over the entire distance traveled (essentially a cosmic highway), so no scattering is expected. As a result, for particles with $\mu=1$, the propagation is nearly ballistic, and the associated delays are negligible. For $10<\mu<$ 1, though, a delay arises due to particle gyration and the total distance traveled can be approximated as $d\approx D\sqrt{\frac{\mu^{2}+1}{\mu^{2}}}$. Assuming an average pitch angle of $\sim\frac{\pi}{4}$, we expect delays of about $t_{d}/t_{b}\sim 0.7$, consistent with the numerical results shown in Figure 1. For $\ell_{\rm c}>D$, these calculations indicate that particle accumulate a delay of about 2 Myr for every Mpc separating us from the source. These delays alone may exceed typical AGN duty cycles, especially for sources on cosmological distance, though Mpc-long magnetic structures may be rare or even never realized.
Region c
In this region $\ell_{\rm c}<\mathcal{R}<D$ and we see that $t_{d}/t_{b}$ is negligible above $\lambda=D$; instead, for sufficiently small values of $\mathcal{R}/D$ and/or large values of $\ell_{\rm c}/D\lesssim 0.1$, UHECRs are effectively scattered on scales much smaller than their Larmor radii, and hence potentially large delays may arise.
Region d
This is the part of the map where $\mathcal{R}<\ell_{\rm c}<D$. Particles diffuse, and time delays are significant; the initial pitch angle is irrelevant and time delays can easily be $t_{d}/t_{b}\gtrsim 10^{3}$. In this regime even the delay of UHECRs produced in circum-Galactic sources ($D\sim 30-300$kpc) would already exceed $>10^{2}$$10^{3}$ Myr. Delays of several Gyr would likely photodisintegrate all of the heavy elements, at odds with Auger observations (e.g., Abreu et al., 2022; Abbasi et al., 2024), as we discuss in §5.1.
Figure 2: Time delay map $t_{d}$(yr) as a function of the coherence length $\ell_{\rm c}$(pc) and the $B$-field $B$(G) that UHECRs with rigidity $\rho_{\rm max}=E/Z=5\times 10^{18}$ probe over different distances as specified. Particles are initialized with $\mu=\cos{\pi/4}$, though the dependence on $\mu$ is minimal. The right panel corresponds to the distance to the nearest jetted AGN, Centaurus A.
4 Astrophysical Implications
The results presented above are general and apply to different parts of the journey of different CR species; we now discuss the implications for UHECRs in their sources and on galactic and extragalactic scales.
4.1 General Considerations
Figure 2 shows the expected delay, $t_{d}$, for particle propagation as a function of coherence length $\ell_{\rm c}$ and magnetic field strength $B$. Let us consider particles with rigidity $\rho_{\rm max}=5\times 10^{18}$ V over distances of 1 Mpc (the typical size of a galaxy cluster) and 3.5 Mpc, which corresponds to the distance to Centaurus A (Cen A), a nearby AGN (e.g., Ferrarese et al., 2007; Majaess, 2010) with indications of an anisotropic UHECR excess above 40 EeV (Abreu et al., 2022). The rigidity $\rho_{\rm max}$ represents the highest rigidity needed to explain the UHECR flux and composition (e.g., Aloisio et al., 2014; Aab et al., 2017), and thus minimizes the possible delay. For comparison, we also show the boundaries where $\lambda=D$ (introduced in §3 and Figure 1) as solid and dashed lines. When the delay becomes comparable with the age of the Universe, an effective magnetic horizon arises (e.g., Parizot, 2004; Globus and Piran, 2017; Globus et al., 2019; Abdul Halim et al., 2024), which has the effects of 1) limiting the observable Universe, 2) introducing a low-energy cut-off in the UHECR spectra, and 3) completely depleting the heavy elements. These last two points may be important to account for the light chemical composition observed at $E\lesssim 10^{18}$eV and the relatively smooth transition between Galactic and extra-galactic CRs (e.g., Aloisio et al., 2012; Unger and Farrar, 2017).
4.2 Delays Due to Propagation in the Source
During the initial propagation phase, particles can spend a significant amount of time in regions inside or immediately around their sources. In the case of AGNs, such regions (e.g., jet cocoons, lobes) can have sizes that reach 100kpc with $\mu$G fields. From Figure 1, we can retrieve the maximum associated delays to be a few Myr. This heavily depends on the size of the regions, but could be another way in which particles are significantly delayed. For instance, tracking of accelerated particles in relativistic MHD simulations confirm that the cocoon is effective in isotropizing the UHECR distribution released by an AGN jet (Mbarek and Caprioli, 2019, 2021).
4.3 Propagation Inside Galaxy Clusters
The left panel of Figure 2 shows the delays expected on a distance of 1 Mpc, the typical size of a galaxy cluster, in particular for the typical cluster values of $B\lesssim 0.1-1\rm\mu G$ and $\ell_{\rm c}\sim 10^{2}-10^{4}\rm pc$ (magenta box); such delays can already be comparable to or greater than the duty cycles of AGNs (see also Berezinsky et al., 1997; Blasi and Olinto, 1998). This is particularly important since powerful jetted AGNs are predominately found in galaxy clusters (e.g., Begelman et al., 1984; Best et al., 2007; Fang and Murase, 2018), and are prominent UHECR source candidates (e.g., Caprioli, 2015; Mbarek and Caprioli, 2019; Katz et al., 2009; Murase and Fukugita, 2019; Jiang et al., 2021). Also, over these timescales, a galaxy could change its identity and flip between AGN, starburst, or quiescent states, thus affecting any study that seeks a correlation with a specific class.
4.4 Centaurus A as a potential UHECR source
[…]
While today Cen A looks like a rather weak FR-I, in the past it could have hosted a more powerful jet able to accelerate particles up to UHECR levels (e.g., Mbarek and Caprioli, 2021; Mbarek et al., 2025). Delays $\gtrsim$ 1 Myr may then imply large uncertainties in the actual efficiency of Cen A at accelerating UHECRs (an argument valid for any source). For instance, a rather large fraction of Cen A’s current bolometric luminosity would need to be converted into energetic particles, so that this AGN contributes a large fraction of the UHECR flux at Earth (Mbarek et al., 2025); this request would be eased if Cen A’s jet were more luminous several Myr ago.
4.5 UHECR Deflections During Transport
Time delays go along with angular deflections in the apparent position of the source, an effect well investigated for the GMF (e.g., Takami et al., 2012; Jansson and Farrar, 2012; Alves Batista et al., 2016; Unger and Farrar, 2017, 2024; Bister et al., 2024). Even at the new of GMF deflections, one can only trace back UHECRs to sources if $\lambda\gg D$ and $\mathcal{R}>\ell_{\rm c}$.
We can calculate trajectory deflection in our framework, and results are consistent with the calculations originally done in the diffusive regime, e.g., by Waxman and Miralda-Escude (1996). If we assume a constant magnetic field magnitude, the average deflection will be $\langle\theta\rangle\sim\frac{\ell_{\rm c}}{2\mathcal{R}}$ and the overall deflection $\theta_{d}$ will be $\theta_{d}\sim\langle\theta\rangle\sqrt{D/\ell_{\rm c}}$ after a distance D.
In particular, we calculate that for a distance $D=1$Mpc and $\rho=\rho_{\rm max}$, practically all particles with $\lambda\leq D$ are deflected by more than 20. As for particles with $\lambda>D$, 10% are deflected by more than 10. This underscores the importance of including deflections due to extragalactic magnetic fields in propagation models when $\lambda\leq D$, an effect that adds on the time delay in hindering the correlation between present-day potential sources and UHECR arrival directions.
Figure 3: Upper Panel: Mean free path $\lambda_{A}$ for one photodisintegration interaction with CMB+COB+CIB due to the GDR for different species. Dashed lines represent the maximum UHECR energy $5Z\times 10^{18}$eV. Dotted lines represent the contribution of the CMB only. Lower Panel: Maximum detectable atomic mass $\langle A_{\rm d}\rangle$ as a function of $d_{\rm tot}$, the total distance traveled for UHECRs with rigidity $\rho_{\rm max}=5\times 10^{18}$eV. The blue shaded area represents CNO-like detected particles ($\langle A_{\rm d}\rangle\in[11,17]$), and the pink shaded area is for He-like particles ($\langle A_{\rm d}\rangle\in[3,5]$).
5 Effects of Losses
Several studies have explored how photodisintegration processes influence the UHECR horizon (Aloisio et al., 2012; Globus et al., 2023; Abdul Halim et al., 2024). In this work, we further these investigations by presenting some generalized constraints on the properties of the extragalactic magnetic fields, i.e., magnitude and coherence lengths. We also assess whether more precise measurements of heavy chemical compositions could refine current constraints.
Figure 4: Map of the Distance traveled as a function of $\ell_{\rm c}$(pc) and $B$(G) that iron (Fe) and carbon (C) UHECRs with rigidity $\rho_{\rm max}=E_{\rm Fe}/Z\simeq 5\times 10^{18}$ probe. The contour plots show the maximum detected Atomic mass A of the parent UHECR (Fe or C) after propagation to Earth because of losses due to the CMB, CIB, and COB.
5.1 General Considerations
The upper panel of Figure 3 shows the mean free path $\lambda_{A}$ for different UHECR species with maximum rigidity $\rho_{\rm max}=5\times 10^{18}$eV. We find our GDR calculations to be consistent with previous studies including Allard (2012), SimProp (e.g., Aloisio et al., 2012), and CRPropa (e.g., Batista et al., 2015; Alves Batista et al., 2022). Ultimately, photodisintegration introduces a maximum detectable atomic mass $\langle A_{\rm d}\rangle$ for each injected heavy species from a particular source, which depends on the total distance traveled, $d_{\rm tot}$, as shown in the lower Panel of Figure 3.
If UHECR sources are distributed across a range of distances, a fraction of the nuclei originating from nearby sources may arrive at Earth with their original atomic mass. As a result, this effect can extend the high-energy cutoff beyond what is expected from a purely distant source distribution, analogous to the energy suppression imposed by the Greisen–Zatsepin–Kuzmin effect (Greisen, 1966; Zatsepin and Kuz’min, 1966). This fraction of intact nuclei may contribute to a softer but detectable high-energy tail in the UHECR spectrum, reflecting the mixed composition and spatial distribution of the sources.
Before discussing our results, it is useful to lay out some preliminary considerations. Since the reconstruction of the UHECR shower properties points to the most energetic UHECRs being as heavy as CNO ($\langle A_{d}\rangle\sim 12$) (Halim et al., 2023), we can devise two scenarios for the origin of such elements based on $d_{\rm tot}$. Either i) detected CNOs are primaries that barely get photodisintegrated, which requires $d_{\rm tot}\leq$100 Mpc (see bottom panel of Figure 3), or ii) they are secondary particles that come from photodisintegrated, Fe-like UHECRs, since Fe is likely the most abundant species among those with $A>A_{\rm CNO}$, as in Galactic CRs (e.g., Hörandel et al., 2006; Caprioli et al., 2010, 2017).
The scenario with secondary CNOs (from primary Fe-like) is quite fine tuned because it would imply i) a dearth of He-like UHECRs, which are instead quite abundant, and ii) that the total distance traveled by most, if not all, Fe-like nuclei must be exactly around $\sim$600-800 Mpc. An injection spectrum rich in Si-like nuclei could ease these requirements, but it is disfavored by standard nucleosynthesis channels. The scenario with primary CNOs, on the other hand, seems more plausible, as the main requirement that needs to be satisfied is to have sufficient sources within $d_{\rm tot}\leq$100 Mpc.
5.2 Centaurus A as a potential UHECR source: Constraints From Losses
In the following, we discuss a scenario where a significant fraction of the UHECR flux stems from Cen A (e.g., Abreu et al., 2022). Assuming that the UHECR injection composition includes CNO and Fe, Figure 4 shows a map of the traveled distance as a function of $B$ and $\ell_{\rm c}$ from a source at $D=3.5$ Mpc, the distance to Cen A. Overlaid on the map are the contours indicating the maximum detectable atomic mass, $\langle A_{\rm d}\rangle$, of Fe and C after propagation to Earth. For reference, we also include the limits defined in Figure 2. We observe that a scenario where UHECR CNOs from Cen A are primary particles is feasible, with $A_{\rm d}(\text{C})\simeq 11$–12 for $\rho_{\rm max}$ if $B\sim 0.1-1\mu$G and $\ell_{\rm c}\leq 10^{4}$ pc, as outlined above. Conversely, a scenario in which UHE CNOs from Cen A are secondaries (with Fe-like parent particles) would require the magnetic fields between us and Cen A to be unrealistically strong, $B\gtrsim 1\mu$G (left panel of Figure 4). Furthermore, we note that the expected atomic mass of the parent Fe particles would be $A_{\rm d}(\text{Fe})\simeq 55-56$, suggesting that detecting UHECRs with atomic masses $A\gtrsim 20$ would be plausible.
6 Conclusions
In this paper we considered the propagation of UHECRs across a wide range of magnetic field strengths and coherence lengths ($\ell_{\rm c}$), examining the ensuing time delay, angular deflection, and attenuation losses as they travel toward Earth.
[…]
In general, large time delays and angular deflections may render it impossible to correlate UHECR arrival directions not just with short-lived sources (gamma-ray burst, tidal disruption events, newly-born ms pulsars), but even with sources with much longer duty cycles, such as AGNs and starbursts galaxies.
Finally, we note how more precise measurements of the UHECR chemical composition at the highest energies, coupled with $B$$\ell_{\rm c}$ delay maps and photodisintegration calculations, could constrain extragalactic magnetic fields, particularly when combined with observations of UHECR anisotropies that hint at potential sources such as AGNs or starburst galaxies (e.g., Halim et al., 2023). For instance, the survival of Fe-like UHECRs from a known distance would put an upper limit on the average strength of extragalactic magnetic fields. Such a detection is a reasonable expectation, given the relative abundance of Fe in Galactic CRs and the observed trend of heavier UHECR composition at higher energies (Abreu et al., 2022).
Acknowledgements
We would like to thank Pasquale Blasi and Kohta Murase for their valuable feedback on the material covered by this work, and two anonymous referees for their comments, which helped improving the manuscript. Simulations were performed on computational resources provided by the University of Chicago Research Computing Center. This work was supported by NSF through grants AST-2009326 and AST-2308021.